OpenStudy (anonymous):

The ratio of the length to the width of a map is 5 to 2. The map is exactly 15 inches in length. What is the width of the map? A) 6 inches B) 7 inches C) 10 inches D) 12 inches

OpenStudy (anonymous):

6 inches

OpenStudy (anonymous):

5/2 = 15/x x= (15)(2/5) x=6